Exercise 7-25 Logan Bruno Company has just received the August 31, 2014, bank statement, which is summarized below County National Bank Balance, August 1 Deposits during August Note collected for depositor, including $98 interest Checks cleared during August Bank service charges Balance, August 31 Disbursements Receipts Balance $22,851 $78,536 101,387 2,537 103,924 19,778 19,729 19,729 $84,146 The general ledger Cash account contained the following entries for the month of August. Balance, August 1 Receipts during August 24,512 Disbursements in August 86,374 85,365 Deposits in transit at August 31 are $9,268, and checks outstanding at August 31 total $3,780. Cash on hand at August 31 is $756. The bookkeeper improperly entered one check in the books at $146.50 which was written for $164.50 for supplies (expense); it cleared the bank during the month of August. Prepare a bank reconciliation dated August 31, 2014, proceeding to a correct balance
Answer #1

Answer

Account Title & Explaination Debit Credit August 31,2014 Cash Account 2537 Note Recievable Interest Revenue (to record collec

Working

Error in Company Cash account = 164.50 - 146.50 = $ 18

Balance as per Book Aug 31 before correction = 24512 + 85365 - 86374 = $ 23503